# A path-dependent PDE solver based on signature kernels

### v3

@@ 2.1. Path spaces
$\mathcal{C}_{t}:=\Big{\{}\gamma\in D(\mathbb{T},\mathbb{R}^{e}):\gamma\!\mid_{[t,T]}\in\mathcal{C}([t,T],\mathbb{R}^{e})\Big{\}}.$
This space allows to consider paths of the form $\gamma+\varepsilon\eta\mathbbm{1}_{[t,T]}$ where $\gamma,\eta\in\mathcal{C}(\mathbb{T},\mathbb{R}^{e})=\mathcal{C}_{0}$ and $\varepsilon>0$ which in turn permits the definition of pathwise derivatives on this space. In parallel, for $p\geq 1$, we also leverage the $p$-variation seminorm which is defined for $0\leq t<t^{\prime}\leq T$ and a path $\gamma:[t,t^{\prime}]\to\mathbb{R}^{e}$ as
$\left\lvert\gamma\right\rvert_{p-var;[t,t^{\prime}]}:=\left(\sup_{(t_{i})_{i}\in\mathcal{D}}\sum_{i}\left\lvert\gamma_{t_{i+1}}-\gamma_{t_{i}}\right\rvert^{p}\right)^{1/p},$
where $\mathcal{D}$ is the set of all partitions of the form $t=t_{0}<t_{1}<\cdots<t_{n}=t^{\prime}$ for some $n\in\mathbb{N}$. The associated norm is $\left\lVert\gamma\right\rVert_{p-var;[t,t^{\prime}]}=\left\lvert\gamma_{t}\right\rvert+\left\lvert\gamma\right\rvert_{p-var;[t,t^{\prime}]}$. This gives rise to the spaces $C^{p-var}([t,t^{\prime}],\mathbb{R}^{e})$ of continuous paths with one strictly monotone coordinate  [^1] This is a technical assumption needed to ensure injectivity of the signature and consequently the point separating property assumption needed to establish universality of the signature kernel by Stone-Weierstrass arguments. More generally one could work with equivalence classes obtained by quotienting $\mathcal{C}$ by the so-called tree-like equivalence relation $x\sim_{\tau}y\iff S(x)=S(y)$; with the assumption that the paths have one strictly monotone coordinate, each equivalence class collapses to a singleton, which is the path itself. The monotone coordinate is usually taken to be time. and finite $p$-variation norm. For each such path $\gamma$ we denote by $\gamma^{0}$ the strictly monotone coordinate and by $\check{\gamma}$ the $\mathbb{R}^{e-1}$-valued path without the strictly monotone coordinate.
The case $p=1$ is called the bounded variation space and also denoted $\mathrm{BV}([t,t^{\prime}],\mathbb{R}^{e})$. Of particular interest to us are the spaces $C^{p-var}_{t}:=C^{p-var}([t,T],\mathbb{R}^{e})$ for some $t\in\mathbb{T}$. While the path space $\mathcal{C}_{t}$ in the original theory of [54] allows for one jump, this would lead to severe complications for the application of signature kernels. We show in the next subsection how to adapt our framework to avoid the introduction of jumps. Furthermore, we want to equip the space $\mathrm{BV}_{0}:=C^{1-var}_{0}$ with a topology that renders the signature map continuous (see Section 3 for the details). We could choose any $1\leq p$-variation topology for this task, however the necessity to characterise compact sets of $\mathrm{BV}_{0}$ later on forces us to consider $p>1$. This task involves the introduction of Hölder spaces $C^{\alpha}([t,t^{\prime}],\mathbb{R}^{e})$, with $\alpha\in(0,1)$ endowed with the $\alpha$-Hölder norm
If the interval $[t,t^{\prime}]$ is clear from the context, for instance if $\gamma\in\mathrm{BV}([t,T],\mathbb{R}^{e})$, we omit to write it as a subscript. Having in mind to characterise compact subsets of $C^{p-var}$, we follow [20] and intersect it with a Hölder space: let $C^{p-var,\alpha}([t,t^{\prime}],\mathbb{R}^{e}):=C^{p-var}([t,t^{\prime}],\mathbb{R}^{e})\cap C^{\alpha}([t,t^{\prime}],\mathbb{R}^{e})$ be the space of $\alpha$-Hölder continuous paths with finite $p$-variation. The corresponding norm is $\left\lVert\gamma\right\rVert_{p-var,\alpha;[t,t^{\prime}]}:=\left\lvert\gamma_{t}\right\rvert+\left\lvert\gamma\right\rvert_{\alpha;[t,t^{\prime}]}+\left\lvert\gamma\right\rvert_{p-var;[t,t^{\prime}]}$. It is then proven in [20, Theorema A.8] that the embedding $C^{p-var,\alpha}\hookrightarrow C^{p^{\prime}-var,\alpha^{\prime}}$ is compact for all $1\leq p<p^{\prime}\leq\infty$ and $10\leq\alpha^{\prime}<\alpha<$ 1 such that $\alpha p<1$ and $\alpha^{\prime}p^{\prime}<1$. In particular, bounded subsets of $C^{1-var,\alpha}$ are compact with respect to the $p$-variation topology for all $\alpha\in(0,1)$ and $p>1$, see also [25, Proposition 5.30].
$\widetilde{\Omega}=\Big{\{}(t,x,\gamma)\in\mathbb{T}\times\mathbb{R}^{d}\times\mathcal{C}_{t}\Big{\}},\qquad\widetilde{\bm{d}}((t,x,\gamma),(\bar{t},\bar{x},\bar{\gamma})):=\left\lvert t-\bar{t}\right\rvert+\left\lvert x-\bar{x}\right\rvert+\left\lVert\gamma-\bar{\gamma}\right\rVert_{0;\mathbb{T}}.$
Note that $\gamma$ lives in $\mathcal{C}_{t}$ which depends directly on $t$. We will denote by $\Omega$ the state space with $\mathrm{BV}_{0}$ instead of $\mathcal{C}_{t}$ and such that paths are constant on $[0,t]$ (except for the strictly monotone coordinate), that is
$\Omega:=\big{\{}(t,x,\gamma)\in\mathbb{T}\times\mathbb{R}^{d}\times\mathrm{BV}_{0}:\check{\gamma}_{s}=\check{\gamma}_{0}\;\text{for all }s\leq t\big{\}}.$
The analogous distance is then
$\bm{d}((t,x,\gamma),(\bar{t},\bar{x},\bar{\gamma})):=\left\lvert t-\bar{t}\right\rvert+\left\lvert x-\bar{x}\right\rvert+\left\lVert\gamma-\bar{\gamma}\right\rVert_{p-var;\mathbb{T}}.$
2.2. Functions on continuous paths
This subsection presents the framework of [54]. Let $\mathcal{C}(\widetilde{\Omega}):=\mathcal{C}(\widetilde{\Omega},\mathbb{R})$ denote the set of real-valued functions $u:\widetilde{\Omega}\to\mathbb{R}$ continuous under $\widetilde{\bm{d}}$. For $u\in\mathcal{C}(\widetilde{\Omega})$, define the right time derivative
@@ 2.2. Functions on continuous paths
for all $t\in[0,T)$, provided the limit exists. By convention we set $\partial_{t}u(T,x,\gamma)=0$. The derivative $\partial_{x}u(t,x,\gamma)$ is defined in the same way for all $(t,x,\gamma)\in\widetilde{\Omega}$. We also introduce the directional derivative ~$\widetilde{\partial}_{\gamma}^{\eta}u(t,x,\gamma)$ with respect to $\gamma\in\mathcal{C}_{0}$, in the direction of $\eta\in\mathcal{C}([t,T],\mathbb{R}^{e})$:
@@
Notice that the path is only perturbed on the interval $[t,T]$, hence the necessity of a path space allowing for jumps. The operator $\eta\mapsto\widetilde{\partial}_{\gamma}^{\eta}u$ is linear on $\mathcal{C}([t,T],\mathbb{R}^{e})$ and a Fréchet derivative; moreover if it exists it is equal to the Gateaux derivative. We observe crucially that both the Fréchet and Gateaux derivative are operators acting on functions from $\widetilde{\Omega}$ to $\mathbb{R}$; in particular the time dependence is essential. We define higher derivatives as linear operators on $(\mathcal{C}([t,T],\mathbb{R}^{e}))^{n}$, in particular the second derivative is a bilinear operator on $\mathcal{C}([t,T],\mathbb{R}^{e})\times\mathcal{C}([t,T],\mathbb{R}^{e})$ and we denote it ~$\widetilde{\partial}_{\gamma}^{\eta{\bar{\eta}}}u(t,x,\omega)$ for $\eta,{\bar{\eta}}\in\mathcal{C}([t,T],\mathbb{R}^{e})$.
Let $n,n_{1},n_{2},n_{3}\in\mathbb{N}_{0}$. We denote the higher derivatives $\partial_{t}^{n_{1}},\,\partial_{x}^{n_{2}}$ and ~$\widetilde{\partial}_{\gamma}^{\bm{\eta}}$, for $\bm{\eta}=(\eta_{1},\cdots,\eta_{n_{3}})\in(\mathcal{C}([t,T],\mathbb{R}^{e}))^{n_{3}}$; the order of the directional derivative is implied by the dimension of $\bm{\eta}$. We say that a function $u\in\mathcal{C}(\widetilde{\Omega})$ belongs to $\mathcal{C}^{\mathcal{N}}(\widetilde{\Omega})$ if the linear operator $\partial_{t}^{n_{1}}\,\partial_{x}^{n_{2}}\,\widetilde{\partial}_{\gamma}^{\bm{\eta}}$ exists and is continuous with respect to $\widetilde{\bm{d}}$ for all $(n_{1},n_{2},n_{3})\in\mathcal{N}\subset\mathbb{N}^{3}_{0}$. The examples of interest arise from Kolmogorov equations, hence only the first order time derivative appears and is never combined with the spatial and pathwise derivatives, while the latter are present up to order two. The corresponding subset is $\mathcal{N}=\{(n_{1},n_{2},n_{3})\in\mathbb{N}^{3}_{0}:n_{1}\leq 1,n_{2}=n_{3}=0,\text{ or }n_{1}=0,n_{2}+n_{3}\leq 2\}$.
@@
(2.2) ~$\widetilde{\partial}_{\gamma}^{K^{t}}u(t,x,\gamma)$ $:=\lim_{\delta\downarrow 0}\widetilde{\partial}_{\gamma}^{K^{\delta,t}}u(t,x,\gamma),$
@@ 2.3. The target PDEs
This paper’s focus is on approximating functionals $u$ which only depend on the path over $[t,T]$, that is there exist $\tilde{u}$ such that $u(t,x,\gamma)=\tilde{u}(t,x,\gamma\lvert_{[t,T]})$. This is the case of all the examples presented in this section. These functionals are solutions to linear PPDEs taking the general form for all $(t,x,\gamma)\in\widetilde{\Omega}$
(2.3) {~$\left\{\begin{array}[]{rl}&\widetilde{\mathcal{L}}_{t}u(t,x,\gamma)=g(t,x,\gamma),\\ &u(T,x,\gamma)=\phi(T,x,\gamma),\end{array}\right.$
where $f,\phi:\widetilde{\Omega}\to\mathbb{R}$ are measurable functions and the linear operator ~$\widetilde{\mathcal{L}}_{t}$ is a combination of derivatives of all arguments. More precisely, there exist $\mathcal{N}\subset\mathbb{N}^{3}_{0}$, $a_{n}\in\mathbb{R}$ for all $n\in\mathcal{N}$ and $\bm{\eta}\in(\mathcal{C}([t,T],\mathbb{R}^{e}))^{n_{3}}$ such that
(2.4) ~$\widetilde{\mathcal{L}}_{t}:=\sum_{n=(n_{1},n_{2},n_{3})\in\mathcal{N}}a_{n}\partial_{t}^{n_{1}}\,\partial_{x}^{n_{2}}\,\widetilde{\partial}_{\gamma}^{\bm{\eta}},$
@@
More precisely, this paper will focus on three examples of parabolic second order PDEs which arise from the Feynman-Kac formula. The solution to these PPDEs thus have a probabilistic representation, a crucial advantage to assess the accuracy of our method. However we emphasise that the numerical scheme does not use this a priori knowledge as in supervised learning tasks. Much like deep Galerkin or PINN algorithms [51, 43], our method can be deployed to solve PPDEs for which no numerical method is known.
Our examples take as underlying a Gaussian Volterra process $\widehat{W}$ defined as $\widehat{W}_{t}:=\int_{0}^{t}K(t,r)\,\mathrm{d}W_{r}$ for some kernel $K\in L^{2}(\mathbb{T}^{2},\mathbb{R})$. The key object, already introduced in the introduction, is the double index process, for $0\leq t\leq s$, $\Theta^{t}_{s}=\mathbb{E}_{s}[\widehat{W}_{t}]=\int_{0}^{t}K(s,r)\,\,\mathrm{d}W_{r}$. In alignment with the prolific rough volatility literature, we will mostly focus on the specification $K(t,r)=(t-r)^{H-1/2}\mathbbm{1}_{r\in[0,t)}$ for $H\in(0,1)$ but other choices are possible, see for instance [10, Assummption 2.6]. The current setup (introduced in [54]), as well as our numerical method, allow for path-dependent payoffs. Since $\widehat{W}$ is a fractional process, the path-dependence is inherent and transfers to the PDE even for state-dependent terminal conditions (payoffs). For clarity of exposition and because well-posedness results are scarcer in the path-dependent case, we will however stick to state-dependent payoffs in our numerical experiments.
In each of the three subsequent subsections, we present (i) the model (ii) the value function $u$ and (iii) the PPDE, in that order. We point to the suitable references for well-posedness and for more details. We emphasise that all three examples are particular cases of [10, Proposition 2.14]; this result ensures, under general assumptions, existence and uniqueness of a classical solution to the PPDE associated to a Volterra process.
@@ 2.3.1. Fractional Brownian motion
(2.5) $u(t,\gamma):=$ $\,\mathbb{E}\left[\phi(\widehat{W}_{T})+\int_{t}^{T}f_{s}(\widehat{W}_{s})\,\mathrm{d}s\Big{|}\Theta^{t}=\gamma\right]$
(2.6) $=$ $\,\mathbb{E}\left[\phi\left(\gamma_{T}+\int_{t}^{T}K(T,s)\,\mathrm{d}W_{s}\right)+\int_{t}^{T}f_{s}\left(\gamma_{s}+\int_{t}^{s}K(t,r)\,\mathrm{d}W_{r}\right)\,\mathrm{d}s\right].$
In particular, we observe that $u(t,\Theta^{t})=\mathbb{E}\left[\phi(\widehat{W}_{T})+\int_{t}^{T}f_{s}(\widehat{W}_{s})\,\mathrm{d}s\Big{|}\mathcal{F}_{t}\right].$ Note that the path-dependence of the conditional expectation is contained in the path $\Theta^{t}$ because the payoff only acts on $\widehat{W}_{s}$ for $s\in[t,T]$, whereas a payoff depending on $\widehat{W}_{s}$ for all $s\in[0,T]$ would require to consider the concatenated path $\widehat{W}_{s}\mathbbm{1}_{s<t}+\Theta^{t}_{s}\mathbbm{1}_{s\geq t}$.
Based on [54, Theorem 4.1], under some standard regularity assumptions on $\phi,\psi$ and $K$, $u$ is a solution to the path-dependent heat equation for all $(t,\gamma)\in\widetilde{\Omega}$
(2.7) $\left\{\begin{array}[]{rl}&\partial_{t}u(t,\gamma)+\frac{1}{2}\widetilde{\partial}^{K^{t}K^{t}}_{\gamma}u(t,\gamma)+\psi(t,\gamma)=0,\\ &u(T,\gamma)=\phi(\gamma_{T}).\end{array}\right.$
It can also be proved [10] that this solution is unique in the space where the functional Itô formula [54, Theorem 3.17] is applicable.
@@ 2.3.2. VIX options under rough Bergomi model
Options on VIX and realised variance form an important class of derivatives, which can be seen as options with path-dependent payoffs. We define ${\rm VIX}_{T}^{2}=\frac{1}{\Delta}\int_{T}^{T+\Delta}\mathbb{E}_{T}[\psi_{s}(\widehat{W}_{s})]\,\mathrm{d}s$ where $\Delta$ is a positive constant corresponding to 30 days, and $\widehat{W}$ is a Gaussian Volterra process in $\mathbb{R}^{e}$ for any $e\in\mathbb{N}$. It is shown in [42] that an option on the VIX (future) has the representation
@@
where $J^{t,T}_{s}=\Theta^{T}_{s}-\Theta^{t}_{s}$ is independent of $\mathcal{F}_{t}$ and $\mathfrak{F}(\gamma):=\frac{1}{\Delta}\int_{T}^{T+\Delta}\mathbb{E}[\psi_{s}(\gamma_{s}+I^{T}_{s})]\,\mathrm{d}s$ where $I^{T}_{s}\sim\mathcal{N}\left(\bm{0}_{d},\int_{0}^{s}K(T,r)K(T,r)^{\top}\,\mathrm{d}r\right)$. Under natural assumptions on $\phi,\psi$ and $K$, the map $u\big{(}t,\gamma):=\mathbb{E}\Big{[}\phi\circ\mathfrak{F}\Big{(}\big{\{}\gamma_{s}+J^{t,T}_{s}:s\in[T,T+\Delta]\big{\}}\Big{)}\Big{]}$ is the unique solution to the PPDE for all $(t,\gamma)\in\widetilde{\Omega}$
(2.8) $\left\{\begin{array}[]{rl}&\displaystyle\partial_{t}u(t,\gamma)+\frac{1}{2}\widetilde{\partial}^{K^{t}K^{t}}_{\gamma}u(t,\gamma)=0,\\ &\displaystyle u(T,\gamma)=\phi\circ\mathfrak{F}(\gamma).\end{array}\right.$

### v4

@@ 2.2. Functions of continuous paths
for all $t\in[0,T)$, provided the limit exists. By convention we set $\partial_{t}u(T,x,\gamma)=0$. The derivative $\partial_{x}u(t,x,\gamma)$ is defined in the same way for all $(t,x,\gamma)\in\widetilde{\Omega}$. We also introduce the directional derivative $\widetilde{\partial}_{\gamma}^{\eta}u(t,x,\gamma)$ with respect to $\gamma\in\mathcal{D}_{0}$, in the direction of $\eta\in\mathcal{C}([t,T],\mathbb{R}^{e})$:
@@
Notice that the path is only perturbed on the interval $[t,T]$, hence the necessity of a path space allowing for a jump at $t$. The operator $\eta\mapsto\widetilde{\partial}_{\gamma}^{\eta}u$ is linear on $\mathcal{C}([t,T],\mathbb{R}^{e})$ and a Fréchet derivative; moreover if it exists it is equal to the Gateaux derivative. We observe crucially that both the Fréchet and Gateaux derivative are operators acting on functions from $\widetilde{\Omega}$ to $\mathbb{R}$; in particular the time dependence is essential. We define higher derivatives as linear operators on $(\mathcal{C}([t,T],\mathbb{R}^{e}))^{n}$, in particular the second derivative is a bilinear operator on $\mathcal{C}([t,T],\mathbb{R}^{e})\times\mathcal{C}([t,T],\mathbb{R}^{e})$ and we denote it $\widetilde{\partial}_{\gamma}^{\eta{\bar{\eta}}}u(t,x,\omega)$ for $\eta,{\bar{\eta}}\in\mathcal{C}([t,T],\mathbb{R}^{e})$.
[…]
@@
(2.5) $\widetilde{\partial}_{\gamma}^{K^{t}}u(t,x,\gamma)$ $:=\lim_{\delta\downarrow 0}\widetilde{\partial}_{\gamma}^{K^{\delta,t}}u(t,x,\gamma),$
@@ 2.3. The target PDEs
This paper’s focus is on approximating functionals $u$ which only depend on the path over $[t,T]$, said differently there exist $\tilde{u}$ such that $u(t,x,\gamma)=\tilde{u}(t,x,\gamma\lvert_{[t,T]})$. This is the case of all the examples presented in this section. These functionals are solutions to linear PPDEs taking the general form for all $(t,x,\gamma)\in\widetilde{\Omega}$
(2.6) $\left\{\hskip-11.38092pt\begin{array}[]{rl}&\widetilde{\mathcal{L}}_{t}u(t,x,\gamma)=g(t,x,\gamma),\\ &u(T,x,\gamma)=\phi(T,x,\gamma),\end{array}\right$.
where $g,\phi:\widetilde{\Omega}\to\mathbb{R}$ are measurable functions and the linear operator $\widetilde{\mathcal{L}}_{t}$ is a combination of derivatives of all arguments. More precisely, there exist $\mathcal{N}\subset\mathbb{N}^{3}_{0}$, $a_{n}\in\mathbb{R}$ for all $n\in\mathcal{N}$ and $\bm{\eta}\in(\mathcal{C}([t,T],\mathbb{R}^{e}))^{n_{3}}$ such that
(2.7) $\widetilde{\mathcal{L}}_{t}:=\sum_{n=(n_{1},n_{2},n_{3})\in\mathcal{N}}a_{n}\partial_{t}^{n_{1}}\,\partial_{x}^{n_{2}}\,\widetilde{\partial}_{\gamma}^{\bm{\eta}},$
@@
More precisely, this paper will focus on three examples of parabolic second order PDEs which arise from the Feynman-Kac formula. The solution to these PPDEs thus have a probabilistic representation, a crucial advantage to assess the accuracy of our method. However we emphasise that the numerical scheme does not use this a-priori knowledge as in supervised learning tasks. Much like deep Galerkin or PINN algorithms [53, 43], our method can be deployed to solve PPDEs for which no numerical method is known.
Our examples take as underlying a Gaussian Volterra process $\widehat{W}$ defined as $\widehat{W}_{t}:=\int_{0}^{t}K(t,r)\,\mathrm{d}W_{r}$ for some kernel $K\in L^{2}(\mathbb{T}^{2},\mathbb{R})$. The key object, already introduced in the introduction, is the double index process, for $0\leq t\leq s$, $\Theta^{t}_{s}=\mathbb{E}[\widehat{W}_{s}\lvert\mathcal{F}_{t}]=\int_{0}^{t}K(s,r)\,\,\mathrm{d}W_{r}$. In alignment with the prolific rough volatility literature, we will mostly focus on the specification $K(t,r)=(t-r)^{H-1/2}\mathbbm{1}_{r\in[0,t)}$ for $H\in(0,1)$ but other choices are possible, see for instance [10, Assumption 2.6]. The current setup (introduced in [56]), as well as our numerical method, allow for path-dependent payoffs. Since $\widehat{W}$ is a fractional process, the path-dependence is inherent and transfers to the PDE even for state-dependent terminal conditions (payoffs). For clarity of exposition and because well-posedness results are scarcer in the path-dependent case, we will however stick to state-dependent payoffs in our numerical experiments.
In each of the three subsequent subsections, we present (i) the model (ii) the value function $u$ and (iii) the PPDE, in that order. We point to the suitable references for well-posedness and for more details. We emphasise that all three examples are particular cases of [10, Proposition 2.14]; this result ensures, under general assumptions, existence and uniqueness of a classical solution to the PPDE associated to a Volterra process. In all three examples, the kernel’s singular behaviour around 0 must be dominated by a power-law $t\mapsto t^{H-\frac{1}{2}},H\in(0,\frac{1}{2})$ and functions are assumed three times differentiable with either exponential (for those applying to a Gaussian variable) or polynomial growth (for the others).
@@ 2.3.1. Fractional Brownian motion
(2.8) $u(t,\gamma):=$ $\,\mathbb{E}\left[\phi(\widehat{W}_{T})+\int_{t}^{T}f(s,\widehat{W}_{s})\,\mathrm{d}s\Big|\Theta^{t}=\gamma\right]$
(2.9) $=$ $\,\mathbb{E}\left[\phi\left(\gamma_{T}+\int_{t}^{T}K(T,s)\,\mathrm{d}W_{s}\right)+\int_{t}^{T}f\left(s,\gamma_{s}+\int_{t}^{s}K(t,r)\,\mathrm{d}W_{r}\right)\,\mathrm{d}s\right].$
In particular, we observe that $u(t,\Theta^{t})=\mathbb{E}\left[\phi(\widehat{W}_{T})+\int_{t}^{T}f(s,\widehat{W}_{s})\,\mathrm{d}s\Big|\mathcal{F}_{t}\right].$ Note that the path-dependence of the conditional expectation is contained in the path $\Theta^{t}$ because the payoff only acts on $\widehat{W}_{s}$ for $s\in[t,T]$.
Based on [56, Theorem 4.1], under some standard regularity assumptions on $\phi,f$ and $K$, $u$ is a solution to the path-dependent heat equation for all $(t,\gamma)\in\widetilde{\Omega}$
(2.10) $\left\{\hskip-11.38092pt\begin{array}[]{rl}&\partial_{t}u(t,\gamma)+\frac{1}{2}\widetilde{\partial}^{K^{t}K^{t}}_{\gamma}u(t,\gamma)+f(t,\gamma)=0,\\ &u(T,\gamma)=\phi(\gamma_{T}).\end{array}\right$.
It can also be proved [10] that this solution is unique in the space where the functional Itô formula [56, Theorem 3.17] is applicable.
@@ 2.3.2. VIX options under rough Bergomi model
Options on VIX and realised variance form an important class of derivatives, which can be seen as options with path-dependent payoffs. We define ${\rm VIX}_{T}^{2}=\frac{1}{\Delta}\int_{T}^{T+\Delta}\mathbb{E}[\psi_{s}(\widehat{W}_{s})\lvert\mathcal{F}_{t}]\,\mathrm{d}s$ where $\Delta$ is a positive constant corresponding to 30 days, and $\widehat{W}$ is a Gaussian Volterra process in $\mathbb{R}^{e}$ for any $e\in\mathbb{N}$. It is shown in [42] that an option on the VIX (future) has the representation
@@
where $J^{t,T}_{s}=\Theta^{T}_{s}-\Theta^{t}_{s}$ is independent of $\mathcal{F}_{t}$ and $\mathfrak{F}(\gamma):=\frac{1}{\Delta}\int_{T}^{T+\Delta}\mathbb{E}[\psi_{s}(\gamma_{s}+I^{T}_{s})]\,\mathrm{d}s$ where $I^{T}_{s}\sim\mathcal{N}\left(\bm{0}_{d},\int_{T}^{s}K(s,r)K(s,r)^{\top}\,\mathrm{d}r\right)$. Under natural assumptions on $\phi,\psi$ and $K$, the map $u\big(t,\gamma):=\mathbb{E}\Big[\phi\circ\mathfrak{F}\Big(\big\{\gamma_{s}+J^{t,T}_{s}:s\in[T,T+\Delta]\big\}\Big)\Big]$ is the unique solution to the PPDE for all $(t,\gamma)\in\widetilde{\Omega}$
(2.11) $\left\{\hskip-11.38092pt\begin{array}[]{rl}&\displaystyle\partial_{t}u(t,\gamma)+\frac{1}{2}\widetilde{\partial}^{K^{t}K^{t}}_{\gamma}u(t,\gamma)=0,\\ &\displaystyle u(T,\gamma)=\phi\circ\mathfrak{F}(\gamma).\end{array}\right$.